OpenStudy (anonymous):

For which equation would x = 5 be a solution? 4x - 7 = 27 3x + 6 = 21 12 - 2x = 13 7 + 6x = 27

OpenStudy (hoblos):

try replacing x=5 in each equation and test the results... in the second choice 3x + 6 = 21 = 3(5) +6 = 15+6=21